OpenStudy (anonymous):

Lickety Split ice cream comes in a cylindrical container with an inside diameter of 6 inches and a height of 10 inches. The company claims to give the customer 25 scoops of ice cream per container, each scoop being a sphere with a 3-inch diameter. How many scoops will each container really hold?

OpenStudy (anonymous):

vol of a cylinder = pi*r^2*h.diameter is 6 inches.So the radius is 3.So the volume of the cylinder is 90pi The volume of a sphere is 4/3 * pi *r^3.r=3/2=1.5.So the volume of the sphere is now 4/3 * pi * 1.5 *1.5 * 1.5 = 4.5pi Number of cones that the cylinder can hold = 90pi/4.5pi = 20 scoops

OpenStudy (anonymous):

i got it thank